OpenStudy (anonymous):

Find the slope between the following points. (0,3) and (8,-4)

OpenStudy (lynfran):

what is -4-3=

OpenStudy (anonymous):

-7

OpenStudy (lynfran):

good what is 8-0=

OpenStudy (anonymous):

8

OpenStudy (lynfran):

good now what is -7/8=

OpenStudy (anonymous):

-7/8

OpenStudy (loser66):

lol

OpenStudy (lynfran):

good thats your slope

OpenStudy (anonymous):

okay

OpenStudy (lynfran):

now remember that the slope is always rise/run...i.e y/x..ok

OpenStudy (anonymous):

okay so the answer ?

OpenStudy (lynfran):